a) A wire loop of resistance 20 Omega and radius 5.0cm is kept in the plane of this paper in a changing uniform magnetic field vec B .The direction of vec B is perpendicular to the plane of the page and points out of it.If the magnitude of the induced current in the loop is 3.0mA determine the rate of change of the magnitude of the magnetic field bar(B) .


1
Expert's answer
2021-06-09T08:06:05-0400

"\varepsilon _i=IR=-\frac{d\Phi}{dt}=\frac{SdB\sin \alpha}{dt}=\frac{\pi r^2 dB}{dt}\implies"

"{dB}=\frac{IR dt}{\pi r^2},"

"dB=\frac{0.003\cdot 20}{3.14\cdot 0.05^2}=7.6~T."